OpenStudy (anonymous):

The equation of a line is x + 5y = 12. What is the y-intercept of the line?

OpenStudy (anonymous):

12/5 5/12 -12 12

OpenStudy (anonymous):

dude. plug in 0 for x

OpenStudy (anonymous):

I honestly don't get any of this.

geerky42 (geerky42):

At y-intercept, x is 0, so you plug in 0 for x, then solve for y. \(x+5y=12\\~\\0+5y=12\\~\\5y=12\)

OpenStudy (anonymous):

so the answer is 5/12?

geerky42 (geerky42):

To solve for y, you would need to divide both sides by 5.

OpenStudy (anonymous):

think about a coordinate plane. the y-intercept is where the line intercepts the y-axis. at any point on the y-axis, the x-coordinate is 0. so the coordinate of the y-intercept is always (0, y-intercept). same for the x-intercept. the coordinate of the x-intercept will be (x-intercept, 0)

OpenStudy (anonymous):

so what's the answer

OpenStudy (triciaal):

the y intercept is the value of the function when you substitute x = 0 and do the necessary calculations. it is where the function crosses the y- axis

OpenStudy (anonymous):

so what would the answer be

OpenStudy (triciaal):

x + 5y = 12 and x = 0 so 0 + 5y = 12 give 5y = 12 5* 1y = 12 to get 1y divide 5*1y by 5 keep the balance divide the other side = 12/5 y = 12/ 5 (0, 12/5)

OpenStudy (anonymous):

Awesome! thanks

OpenStudy (triciaal):

you are welcome.